Grade 12Physical Chemistry

If 50 mL of 0.2 M KOH is added to 40 mL of 0.5 M
HCOOH, the pH of the resulting solution is
(Ka
= 1.8 × 10–4) :

amogh joshi
11 Years agoGrade 12

2 Answers

Sunil Kumar FP
11 Years ago
KOH + HCOOH ---------HCOOK +H2O
.01 ,02 0
0 .01 .01

pH =pKA +log (salt)/(acid)
=-logK +log (.01/.01)
=3.744
Rishi Sharma
5 Years ago
Dear Student,
Please find below the solution to your problem.

KOH + HCOOH ---------HCOOK +H2O
.01 ,02 0
0 .01 .01
pH =pKA +log (salt)/(acid)
= -logK +log (.01/.01)
= 3.744

Thanks and Regards